OpenStudy (anonymous):

simplified form of (15x^5/24y^8) / (4x^2/8y^4)

OpenStudy (danjs):

exponents, exponents, more more more

OpenStudy (watchmath):

\(\frac{15x^5\cdot 8y^4}{24y^8\cdot 4x^2}\)

OpenStudy (danjs):

multiply by the inverse of the bottom fraction

OpenStudy (danjs):

lol nevermind

OpenStudy (anonymous):

its 15x^5 / 24y^8 / 4x^2/8y^4

OpenStudy (anonymous):

:(

OpenStudy (danjs):

Yeah watchman wrote it correct above

OpenStudy (danjs):

flip the bottom fraction over and multiply them together

OpenStudy (danjs):

recall: \[\frac{ 1 }{ a^n }=\frac{ a ^{-n} }{ 1 }\] or the other way around

OpenStudy (danjs):

\[\frac{ 15*x ^{5-2} }{ 24 *y ^{8-4} }\]

OpenStudy (anonymous):

Make sure to multiply the coefficients from the bottom fraction. Its 15*8 and 24*4